OpenStudy (anonymous):

Explain how you would find the exact value of the square root of 2 over the square root of 8.

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{\sqrt{8}}=\sqrt{\frac{2}{8}}=\sqrt{\frac{1}{4}}=\frac{1}{2}\] if you still wonder

OpenStudy (anonymous):

@HubbaHubba